PERTURBATIONS, MIXINGS, AND SPECTRAL CONSTANTS AS A FUNCTION OF VIBRATIONAL STATE

REBECCA A. H. BUTLER, California Institute of Technology, Jet Propulsion Laboratory, 4800 Oak Grove Dr, Pasadena, CA 91109-8099; DOUGLAS T. PETKIE, Department of Physics, Wright State University, Dayton OH 45435; PAUL HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; \underline{FRANK C. DE LUCIA}, Department of Physics, The Ohio State University, Columbus, OH 43210.

We have observed and assigned spectra in more than 30 excited states of HNO_{3} and ClONO_{2}. While one of the primary motivations of this work has been our interest in the relation between the pure rotational and rotation-vibration spectrum of each of these molecules, the results of these analyses provide a perhaps unique opportunity to explore the relations among the many rotation and distortion parameters that describe these spectra. We will show that although the constants are strongly influenced by the mixings of the states, these perturbations can be removed and constants more directly related to molecular properties obtained.